# The Properties of Little Red Dot Galaxies in the ASTRID Simulation

### v2

@@ 2.2.2 Stellar birth cloud modeling
In addition to the dust model for the ISM described in section 2.2 (Wilkins17; Lachance_24), we implement a stellar birth cloud and nebular emission model to our imaging process for this work. We make this addition because one of the leading proposals for star-dominated LRDs is compact starburst galaxies with very young star populations, where emission from birth clouds could be important (Perez_Gonzalez_2024).
For our nebular emission implementation, we use the built-in nebular emission functionality of SynthObs (wilkins_2020). Similarly, for the birth cloud occlusion, we follow the prescription from Vijayan_2021. This entails calculating the birth cloud optical depth
@@ 2.4 Subpopulations of Interest
In addition to the galaxies which fit the LRD criteria, we also identify two other subpopulations of galaxies within Astrid. First, the population of galaxies which host black holes accreting at or above 50% the Eddington rate (high Eddington black holes), as they are one of the primary proposed sources of AGN-dominated LRDs (Rusakov_2025; Naidu_2025; Durodola_2024). Second, the population of galaxies that are compact and low stellar mass. We use the stellar half mass radius ($\rm r_{half,*}$) calculated by SUBFIND as our indicator of compactness, with a criteria of $\rm r_{half,*}\leq 650pc$. We set the maximum stellar mass to $\rm log(M_{*}/\,M_{\odot})=9.5$. We highlight this population as it should contain the primary candidates for the dimmer LRDs not found in Astrid.
@@ 3 Little Red Dot Identification
We use this mock observation process to create images of all of the subhalos at redshifts 5, 6, 7 and 8 with a subfind stellar masses above $10^{8}\,M_{\odot}$. We use these observations to identify which galaxies fit the criteria of LRDs, and compare them to the LRDs seen in observations. We use criteria comprised of a combination of JWST colors, a minimum magnitude threshold, and a compactness measure, similar to Kocevski_2024.
@@
This criteria is based on the color selection in Perez_Gonzalez_2024. The magnitude threshold on F [^4] W is necessary to remove any very dim galaxies that may fit the other two criteria due to noise. The $\rm F277W-F444W>1$ criterion picks out galaxies that are significantly red in the longer wavelength bands of JWST which is necessary to have the red appearance that is core to LRDs. LRDs are not consistently red across the JWST spectrum, but instead have a more neutral color in the shorter wavelength bands. This is indicated by the $\rm F150W-F200W<0.5$ criteria as it removes galaxies that have a strong red color across both short and long wavelength JWST filters.
While we focus our analysis on the galaxies which fulfill all of the color and magnitude criteria, we also highlight galaxies which fulfill the minimum brightness, and redness criteria (along with the compactness criterion discussed below), but do not fulfill the $\rm F150W-F200W<0.5$ criteria. These galaxies can provide both a comparison point for the LRD galaxies, potentially identifying the source of the change in slope between the shorter and longer wavelengths, and insight into how important this feature is to LRDs, as some works like Akins_2024 remove this criterion entirely. We refer to these galaxies as “no-break LRDs” as they feature the red appearance of LRDs, but do not have the break in their spectra between the short and long wavelength bands.
@@
In addition to the color criteria for LRDs, we use a size criterion in order to ensure the galaxies have the point-like appearance that is central to LRDs. We take a simplified version of the Sérsic effective radius criteria (e.g. Kocevski_2024), which requires that an LRD have a single maximum radius threshold, $\rm r_{e,F444W}\leq 4.5$ pixels. This compactness cut removes 7 of the 24 galaxies which fit the color and magnitude criteria, leaving seventeen total LRDs in the Astrid simulation between redshifts 5 and 8. The exact $r_{e}$ value for each LRD can be found in their mock images in figure 1. When converted to physical units, the median $r_{e}$ is $611$ pc.
@@
The low number of LRDs present in Astrid relative to observations further supports the fact that we are only capturing a small fraction of the observed LRD population. Based on the proposed density of LRDs of $\sim 10^{-4}$ $\rm Mpc^{-3}$ from observations (Matthee_2023; Perez_Gonzalez_2024; Kocevski_2024; Kokorev_2024; Finkelstein_2025), for the volume we are analyzing we would expect to find thousands of LRDs, but we only find seventeen (a density of $\sim 10^{-7}$ $\rm Mpc^{-3}$). This is due to a combination of the missing subpopulation(s), and an under-density of $\sim$2 orders of magnitude in the bright LRD population (F [^4] W $\leq 25.5$) The mass and luminosity functions of the galaxies and black holes in this redshift range in Astrid are largely consistent with observations astrid_galaxy_formation; astrid_BHs, indicating this deficit in LRDs is not the result of an under-density of galaxies or black holes in general. We investigate why more galaxies in Astrid, especially dimmer galaxies, are not LRDs in Section 5 and discuss further in Section 6.1.
The redshift distribution of LRDs in Astrid (redshifts 5, 6, 7, and 8 contain 7, 4, 5, and 1 LRDs respectively) is roughly consistent with observations, which find the highest LRD density occurs around $\rm z\sim 5-6$ (Kocevski_2024; Pacucci_2025). Due to the low number of LRDs in Astrid it is difficult to make a quantitative comparison with the observational distribution.
Figure 4: Galaxy black hole mass plotted against galaxy stellar mass. The underlying 2D histogram shows the population of all galaxies in Astrid colored by the average F277W-F [^4] W color of galaxies in that pixel, while the red and green points indicate LRDs and no-break LRDs found in Astrid and in JWST observations. The contours contain $1\sigma$ and $2\sigma$ of galaxies in their subpopulation. The purple dataset is comprised of AGN fits from Maiolino_2024_JADES; Harikane_2023; Ubler_2023; Stone_2024; Furtak_2024; Yue_2024_eiger compiled in Pacucci_Loeb_2024. The blue star dataset was produced in Ananna_2024 from the X-ray observations of some LRDs.
@@ 4.1 Galaxy masses
The $\rm M_{BH}-M_{*}$ relation is of particular interest due to recent JWST observations of many black holes which appear over-massive compared to their host galaxies (Pacucci_2023). We compare our LRD sample with two samples of similar galaxies from JWST observations; first, a sample of galaxies with over-massive black holes, many of which reside in LRDs, aggregated in Pacucci_Loeb_2024 (over-massive BH sample; shown in purple in Figure 4), and second, a sample from Ananna_2024 (X-ray sample; shown in blue) which proposes upper bounds on the black hole mass based on the X-ray non-detection of LRDs. While these two samples are not fitting the exact same sources, they illustrate how disparate the proposed stellar and black hole masses for galaxies with similar properties can be, depending on the observational features that are prioritized, and the models used during fitting. This is further supported by the potential for bias in host stellar mass fits found by Berger_2025. The dataset compiled in Pacucci_Loeb_2024 suggests over-massive black holes, most of which cannot be reproduced in Astrid, are present in the LRD population. The dataset from Ananna_2024 has generally lower black holes masses, and supports star-dominated, or hybrid star+AGN LRDs.
@@ 4.2.1 Band-specific contribution and attenuation
The bottom two panels of Figure 5 show the attenuation of both the stellar and AGN components of both the overall galaxy population and the LRDs in Astrid. In the F [^4] W band, nearly all LRDs (with the exception of the no-break LRD subpopulation discussed above) have attenuation values ranging from 1 to 4 for both their AGN and stellar components which is consistent with fitted attenuation values from observations (Kokorev_2024; Ma_2024; Akins_2024). In the F150W band, LRDs have AGN attenuation values of 8 to 60, while the stellar attenuation ranges from 2 to 6. This extreme level of AGN attenuation in the shorter wavelengths aligns with attenuation fits from observation (e.g. Brooks_2024). This indicates that the increased dust slope relative to wavelength we employ for AGN attenuation is consistent with observations.
@@ 4.2.2 Galaxy Spectra
The top two panels of Figure 6 show the spectra of galaxies that have similar stellar and black hole masses and luminosities to our LRD population, but fail to appear as LRDs due to their dust content. The top left galaxy has too much dust attenuating both the stars and AGN. This results in a consistent shallow red slope across the spectrum. The top right galaxy is observed along a line-of-sight with minimal attenuation on the AGN. The AGN spectrum is only slightly reddened, and contributes significantly across the JWST spectrum. This indicates how sensitive the appearance of this population of galaxies is to their dust properties. See, in this regard, the model by Ferrara_2023.
@@ 4.3 Stellar and Black Hole Properties
We show the star formation rate (SFR) and stellar half mass radius ($\rm r_{half,*}$) in Figure 7. With the exception of two low mass LRDs, the LRD galaxies have SFRs ranging from typical to well below average for their mass. The bias towards lower SFR among our LRDs suggests they are transitioning towards being quenched likely as a result of feedback from their bright AGN. (Pacucci_Loeb_2024; Pacucci_2024). Within the LRD population, we see the no-break LRDs tend to have lower SFR, which indicates they have been quenched for longer. This is consistent with their redder short-wavelength emission, as their population of young blue stars is smaller.
@@
We also analyze the properties of the black holes present in the galaxies of Astrid. In the left panel of Figure 8, we show the bolometric luminosity of the black holes and compare the overall population to the LRDs. Nearly all of the black holes in the mass range where we find LRDs are accreting at a rate of at least $10\%$ Eddington. This is expected as significant accretion over a long period is the main pathway to producing these high mass black holes. The luminosity of the black hole population in this region is consistent with X-ray and UV observational results (astrid_BHs).
@@ 4.4 Detectability of the AGN in the X-ray band
In addition to the observed JWST colors and sizes, one of the hallmark properties of LRDs is that they are generally undetected in the X-rays, even with very deep stacking analysis (see, e.g., Ananna_2024; Yue_2024; Maiolino_2024; Pacucci_Narayan_2024; Lambrides_2024; Madau_Haardt_2024). This is, in principle, in tension with the hypothesis that the LRDs contain bright AGN.
@@
Given the heavy obscuration necessary for an AGN spectrum to have the steep red slope in the JWST bands, it is suggested that the lack of X-ray detection may be due to the X-ray emission being dimmed below the detectable range by the same gas and dust that is attenuating the visible light (Maiolino_2024). The X-ray emission of the AGN would be attenuated by both the gas in its local environment, and the gas in the ISM along the line-of-sight. The spacial resolution required to fully simulate the AGN environment is well below what is achievable with modern cosmological simulations. As a result we must model this region after the fact via software like CLOUDY as described in section 2.2.1. We calculate the surface density of intervening ISM gas along a line-of-sight from the AGN based on the gas particles of the simulation. This provides a lower bound on the hydrogen surface density that would determine the amount of X-ray attenuation of the AGN, as it does not account for the AGN’s local environment. We calculate the hydrogen column density along 50 lines of sight for each black hole in our population, including the line of sight used for our observations, to provide a systematic view of the ISM component of X-ray attenuation of AGN in our LRDs. We show these results in Figure 9. We see that the bulk of the LRD population falls in the $log(N_{H})=23.0-23.5$ range, while the no-break LRD population extends down to $log(N_{H})=22.5$. This is in alignment with the $N_{H}$ predicted by the Compton thin absorption model proposed in Akins_2024, which employs an AGN SED that is weaker in the X-ray region, due to a low $\alpha_{OX}$ ($\alpha_{OX}=-2.0$) (Pacucci_Narayan_2024). Given the $\rm N(H)$ values we calculate are lower bounds from the ISM only, if we were able to include a self-consistent AGN environment in our calculations, it may raise the $\rm N(H)$ enough such that our results would be consistent with observed X-ray detection limits without relying on an intrinsic X-ray weak AGN spectrum.
@@
Additionally, as discussed in section 2.2.1 we adopt a relatively standard X-ray AGN spectrum, with $\alpha_{OX}=-1.4$ and find our LRDs have intrinsic X-ray luminosities in the range $L_{X}=10^{43}-10^{44}\rm erg/s$ which falls in the region of the upper limits placed on the X-ray luminosity of LRDs should they be AGN-dominated (Ananna_2024). Given our population of LRDs is most comparable to the brightest observed LRDs, it is unsurprising that our LRDs have X-ray luminosities near the upper limits from observations. This would also allow for lower mass, dimmer LRDs with similar overall structure to those found in Astrid to fall below the upper limits placed on AGN $L_{X}$ from observational nondetection.
@@ 5 Where Are The Dimmer LRDs?
The population of LRDs in Astrid represents a subset of the observed population, as their number density is significantly lower than those found in observations, and they are concentrated within the brighter portion of the observed LRD population. Astrid contains $\sim 1/100$ of the expected number of LRDs with UV magnitudes brighter than $\rm M_{UV}=-19$, but completely lacks LRDs with $\rm M_{UV}\geq-18.5$ resulting in an overall LRD number density of $\sim 1/1000$ of observations (Kocevski_2024; Kokorev_2024; Perez_Gonzalez_2024).
@@
This galaxy has an F277W-F [^4] W color of 0.88, which is insufficient to host an LRD. The primary differences between the spectrum of this galaxy and that of the example LRD in Figure 2 are the relative brightness of the AGN and the dust attenuation of both components. In the LRD spectrum, before dust attenuation, the AGN and stellar components have similar brightness in the F [^4] W band. Despite having one of the brightest AGN relative to its stellar mass among the lower stellar mass galaxies, the AGN emission of the dim LRD candidate galaxy is only $\sim 16\%$ as bright as its host galaxy in the F [^4] W band before dust attenuation. This indicates very few, if any, of the lower mass galaxies in Astrid have AGN that are bright enough to significantly impact their overall color. This is likely the result of both black hole accretion and star formation being induced by the presence of cold dense gas, producing host galaxies which are brighter due to their increased stellar mass, and young stellar population. This coupling is broken once AGN feedback begins quenching the host galaxy, as we see in our LRDs and models of high redshift galaxies hosting over-massive black holes (Pacucci_Loeb_2024; Pacucci_2024).
@@
These two factors are heavily impacted by modeling choices in our mock observation pipeline. AGN bolometric luminosity is determined by the simulation, but the AGN SED can impact how that luminosity is distributed, and is dependent on the properties of the gas very near the black hole, which is not resolved by cosmological simulations. Recent work (e.g., Naidu_2025), suggests that AGN embedded in a denser gas cloud, similar to models of direct collapse black holes (e.g. Pacucci_2015; Pacucci_2016), may be a strong candidate for LRD emission, which produces a significantly different SED than our more traditional model. Similarly, our dust model has a normalization factor calibrated in astrid_galaxy_formation, but may not be tuned appropriately for LRD galaxies.
We experiment with these two modeling choices on our dim LRD candidate galaxy. We increase the hydrogen column density of our AGN SED model in cloudy from $10^{23}$ $\rm N/cm^{2}$ to $10^{24}$ $\rm N/cm^{2}$ to approximate the effect of the ”gas-enshrouded" AGN models proposed in works like Inayoshi_2025; Naidu_2025 and, Taylor_2025. Additionally, we increase the dust normalization by a factor of four (from $\kappa_{\rm ISM}=10^{4.1}$ to $\kappa_{\rm ISM}=10^{4.7}$) for the stellar emission to explore the impact of errors in normalization. This also serves as an approximate test of the impact of potential under-density due to simulation effects like resolution limitations and gravitational softening.
@@
The “gas-enshrouded" AGN SED introduces another reddening mechanism to these sources, which could also address the lack of very red LRDs in Astrid. Given the “gas-enshrouded" AGN model we adopted in this experiment was fairly conservative (Naidu_2025 found a Hydrogen column density of $\rm N(H)/cm^{2}\sim 10^{26}$) it is likely that a more extreme version of this model would have a significantly redder spectrum. If such an SED were applied to Eddington or Super-Eddington sources, they may appear similar to the reddest objects seen in observations. We intend to explore these alternative AGN models in more depth in future work.
The increased hydrogen column density in the AGN SED introduces both a steeper intrinsic AGN SED and a significant Balmer break, both of which contribute to the red appearance of this galaxy. These qualities are especially important for any LRDs with minimal dust content, like those suggested by Setton_2025; Xiao_2025, and would result in the consistent Balmer break transition that is found in Setton_2024.
The increased stellar dust attenuation has two primary effects: (I) it reddens the stellar spectrum and (II) it decreases the brightness of the stellar component, which increases the relative contribution of the AGN to the total emission of the source. This increase in AGN contribution is particularly impactful given the significant red color of the “gas-enshrouded" AGN SED. If these high Eddington black holes were present in dimmer, lower mass host galaxies, that would also produce sources with heightened AGN contributions. With the black hole models used in Astrid this is unlikely to occur due to the black hole seeding and feedback models, but it may be possible under alternative models. The BRAHMA simulation suite (Bhowmick_2024) suggests that the over-massive black holes seen in observations are reproducible in simulation with specific seeding techniques, and that much of their mass is the result of black hole mergers. This would allow for the formation of more massive black holes in lower mass galaxies, as their mass growth would be less coupled to star formation. These more massive black holes would be capable of brighter AGN emission without exceeding the Eddington limit, potentially appearing as LRDs.

### v3

@@ 2.2.2 Stellar birth cloud modeling
In addition to the dust model for the ISM described in section 2.2 (Wilkins et al., 2017; LaChance et al., 2024), we implement a stellar birth cloud and nebular emission model to our imaging process for this work. We make this addition because one of the leading proposals for star-dominated LRDs is compact starburst galaxies with very young star populations, where emission from birth clouds could be important (Pérez-González et al., 2024).
For our nebular emission implementation, we use the built-in nebular emission functionality of SynthObs (Wilkins et al., 2020). Similarly, for the birth cloud occlusion, we follow the prescription from Vijayan et al. (2021). This entails calculating the birth cloud optical depth
@@ 2.4 Subpopulations of Interest
In addition to the galaxies which fit the LRD criteria, we also identify two other subpopulations of galaxies within Astrid. First, the population of galaxies which host black holes accreting at or above 50% the Eddington rate (high Eddington black holes), as they are one of the primary proposed sources of AGN-dominated LRDs (Rusakov et al., 2025; Naidu et al., 2025; Durodola et al., 2025). Second, the population of galaxies that are compact and low stellar mass. We use the stellar half mass radius ($\rm r_{half,*}$) calculated by SUBFIND as our indicator of compactness, with a criteria of $\rm r_{half,*}\leq 650pc$. We set the maximum stellar mass to $\rm log(M_{*}/\,M_{\odot})=9.5$. We highlight this population as it should contain the primary candidates for the dimmer LRDs not found in Astrid.
@@ 3 Little Red Dot Identification
We use this mock observation process to create images of all of the subhalos at redshifts 5, 6, 7 and 8 with a subfind stellar masses above $10^{8}\,M_{\odot}$. We use these observations to identify which galaxies fit the criteria of LRDs, and compare them to the LRDs seen in observations. We use criteria comprised of a combination of JWST colors, a minimum magnitude threshold, and a compactness measure, similar to Kocevski et al. (2024).
@@
This criteria is based on the color selection in Pérez-González et al. (2024). The magnitude threshold on F [^4] W is necessary to remove any very dim galaxies that may fit the other two criteria due to noise. The $\rm F277W-F444W>1$ criterion picks out galaxies that are significantly red in the longer wavelength bands of JWST which is necessary to have the red appearance that is core to LRDs. LRDs are not consistently red across the JWST spectrum, but instead have a more neutral color in the shorter wavelength bands. This is indicated by the $\rm F150W-F200W<0.5$ criteria as it removes galaxies that have a strong red color across both short and long wavelength JWST filters.
While we focus our analysis on the galaxies which fulfill all of the color and magnitude criteria, we also highlight galaxies which fulfill the minimum brightness, and redness criteria (along with the compactness criterion discussed below), but do not fulfill the $\rm F150W-F200W<0.5$ criteria. These galaxies can provide both a comparison point for the LRD galaxies, potentially identifying the source of the change in slope between the shorter and longer wavelengths, and insight into how important this feature is to LRDs, as some works like Akins et al. (2024) remove this criterion entirely. We refer to these galaxies as “no-break LRDs” as they feature the red appearance of LRDs, but do not have the break in their spectra between the short and long wavelength bands.
@@
In addition to the color criteria for LRDs, we use a size criterion in order to ensure the galaxies have the point-like appearance that is central to LRDs. We take a simplified version of the Sérsic effective radius criteria (e.g. Kocevski et al., 2024), which requires that an LRD have a single maximum radius threshold, $\rm r_{e,F444W}\leq 4.5$ pixels. This compactness cut removes 7 of the 24 galaxies which fit the color and magnitude criteria, leaving seventeen total LRDs in the Astrid simulation between redshifts 5 and 8. The exact $r_{e}$ value for each LRD can be found in their mock images in figure 1. When converted to physical units, the median $r_{e}$ is $611$ pc.
@@
The low number of LRDs present in Astrid relative to observations further supports the fact that we are only capturing a small fraction of the observed LRD population. Based on the proposed density of LRDs of $\sim 10^{-4}$ $\rm Mpc^{-3}$ from observations (Matthee et al., 2023; Pérez-González et al., 2024; Kocevski et al., 2024; Kokorev et al., 2024; Finkelstein et al., 2025), for the volume we are analyzing we would expect to find thousands of LRDs, but we only find seventeen (a density of $\sim 10^{-7}$ $\rm Mpc^{-3}$). This is due to a combination of the missing subpopulation(s), and an under-density of $\sim$2 orders of magnitude in the bright LRD population (F [^4] W $\leq 25.5$) The mass and luminosity functions of the galaxies and black holes in this redshift range in Astrid are largely consistent with observations Bird et al. (2022); Ni et al. (2022), indicating this deficit in LRDs is not the result of an under-density of galaxies or black holes in general. We investigate why more galaxies in Astrid, especially dimmer galaxies, are not LRDs in Section 5 and discuss further in Section 6.1.
The redshift distribution of LRDs in Astrid (redshifts 5, 6, 7, and 8 contain 7, 4, 5, and 1 LRDs respectively) is roughly consistent with observations, which find the highest LRD density occurs around $\rm z\sim 5-6$ (Kocevski et al., 2024; Pacucci and Loeb, 2025). Due to the low number of LRDs in Astrid it is difficult to make a quantitative comparison with the observational distribution.
Figure 4: Galaxy black hole mass plotted against galaxy stellar mass. The underlying 2D histogram shows the population of all galaxies in Astrid colored by the average F277W-F [^4] W color of galaxies in that pixel, while the red and green points indicate LRDs and no-break LRDs found in Astrid and in JWST observations. The contours contain $1\sigma$ and $2\sigma$ of galaxies in their subpopulation. The purple dataset is comprised of AGN fits from Maiolino et al. (2024b); Harikane et al. (2023); Übler et al. (2023); Stone et al. (2024); Furtak et al. (2024); Yue et al. (2024b) compiled in Pacucci and Loeb (2024). The blue star dataset was produced in Ananna et al. (2024) from the X-ray observations of some LRDs.
@@ 4.1 Galaxy masses
The $\rm M_{BH}-M_{*}$ relation is of particular interest due to recent JWST observations of many black holes which appear over-massive compared to their host galaxies (Pacucci et al., 2023). We compare our LRD sample with two samples of similar galaxies from JWST observations; first, a sample of galaxies with over-massive black holes, many of which reside in LRDs, aggregated in Pacucci and Loeb (2024) (over-massive BH sample; shown in purple in Figure 4), and second, a sample from Ananna et al. (2024) (X-ray sample; shown in blue) which proposes upper bounds on the black hole mass based on the X-ray non-detection of LRDs. While these two samples are not fitting the exact same sources, they illustrate how disparate the proposed stellar and black hole masses for galaxies with similar properties can be, depending on the observational features that are prioritized, and the models used during fitting. This is further supported by the potential for bias in host stellar mass fits found by Berger et al. (2025). The dataset compiled in Pacucci and Loeb (2024) suggests over-massive black holes, most of which cannot be reproduced in Astrid, are present in the LRD population. The dataset from Ananna et al. (2024) has generally lower black holes masses, and supports star-dominated, or hybrid star+AGN LRDs.
@@ 4.2.1 Band-specific contribution and attenuation
The bottom two panels of Figure 5 show the attenuation of both the stellar and AGN components of both the overall galaxy population and the LRDs in Astrid. In the F [^4] W band, nearly all LRDs (with the exception of the no-break LRD subpopulation discussed above) have attenuation values ranging from 1 to 4 for both their AGN and stellar components which is consistent with fitted attenuation values from observations (Kokorev et al., 2024; Ma et al., 2024; Akins et al., 2024). In the F150W band, LRDs have AGN attenuation values of 8 to 60, while the stellar attenuation ranges from 2 to 6. This extreme level of AGN attenuation in the shorter wavelengths aligns with attenuation fits from observation (e.g. Brooks et al., 2024). This indicates that the increased dust slope relative to wavelength we employ for AGN attenuation is consistent with observations.
@@ 4.2.2 Galaxy Spectra
The top two panels of Figure 6 show the spectra of galaxies that have similar stellar and black hole masses and luminosities to our LRD population, but fail to appear as LRDs due to their dust content. The top left galaxy has too much dust attenuating both the stars and AGN. This results in a consistent shallow red slope across the spectrum. The top right galaxy is observed along a line-of-sight with minimal attenuation on the AGN. The AGN spectrum is only slightly reddened, and contributes significantly across the JWST spectrum. This indicates how sensitive the appearance of this population of galaxies is to their dust properties. See, in this regard, the model by Ferrara et al. (2023).
@@ 4.3 Stellar and Black Hole Properties
We show the star formation rate (SFR) and stellar half mass radius ($\rm r_{half,*}$) in Figure 7. With the exception of two low mass LRDs, the LRD galaxies have SFRs ranging from typical to well below average for their mass. The bias towards lower SFR among our LRDs suggests they are transitioning towards being quenched likely as a result of feedback from their bright AGN. (Pacucci and Loeb, 2024; Pacucci et al., 2024). Within the LRD population, we see the no-break LRDs tend to have lower SFR, which indicates they have been quenched for longer. This is consistent with their redder short-wavelength emission, as their population of young blue stars is smaller.
@@
We also analyze the properties of the black holes present in the galaxies of Astrid. In the left panel of Figure 8, we show the bolometric luminosity of the black holes and compare the overall population to the LRDs. Nearly all of the black holes in the mass range where we find LRDs are accreting at a rate of at least $10\%$ Eddington. This is expected as significant accretion over a long period is the main pathway to producing these high mass black holes. The luminosity of the black hole population in this region is consistent with X-ray and UV observational results (Ni et al., 2022).
@@ 4.4 Detectability of the AGN in the X-ray band
In addition to the observed JWST colors and sizes, one of the hallmark properties of LRDs is that they are generally undetected in the X-rays, even with very deep stacking analysis (see, e.g., Ananna et al., 2024; Yue et al., 2024a; Maiolino et al., 2024a; Pacucci and Narayan, 2024; Lambrides et al., 2024; Madau and Haardt, 2024). This is, in principle, in tension with the hypothesis that the LRDs contain bright AGN.
@@
Given the heavy obscuration necessary for an AGN spectrum to have the steep red slope in the JWST bands, it is suggested that the lack of X-ray detection may be due to the X-ray emission being dimmed below the detectable range by the same gas and dust that is attenuating the visible light (Maiolino et al., 2024a). The X-ray emission of the AGN would be attenuated by both the gas in its local environment, and the gas in the ISM along the line-of-sight. The spacial resolution required to fully simulate the AGN environment is well below what is achievable with modern cosmological simulations. As a result we must model this region after the fact via software like CLOUDY as described in section 2.2.1. We calculate the surface density of intervening ISM gas along a line-of-sight from the AGN based on the gas particles of the simulation. This provides a lower bound on the hydrogen surface density that would determine the amount of X-ray attenuation of the AGN, as it does not account for the AGN’s local environment. We calculate the hydrogen column density along 50 lines of sight for each black hole in our population, including the line of sight used for our observations, to provide a systematic view of the ISM component of X-ray attenuation of AGN in our LRDs. We show these results in Figure 9. We see that the bulk of the LRD population falls in the $log(N_{H})=23.0-23.5$ range, while the no-break LRD population extends down to $log(N_{H})=22.5$. This is in alignment with the $N_{H}$ predicted by the Compton thin absorption model proposed in Akins et al. (2024), which employs an AGN SED that is weaker in the X-ray region, due to a low $\alpha_{OX}$ ($\alpha_{OX}=-2.0$) (Pacucci and Narayan, 2024). Given the $\rm N(H)$ values we calculate are lower bounds from the ISM only, if we were able to include a self-consistent AGN environment in our calculations, it may raise the $\rm N(H)$ enough such that our results would be consistent with observed X-ray detection limits without relying on an intrinsic X-ray weak AGN spectrum.
@@
Additionally, as discussed in section 2.2.1 we adopt a relatively standard X-ray AGN spectrum, with $\alpha_{OX}=-1.4$ and find our LRDs have intrinsic X-ray luminosities in the range $L_{X}=10^{43}-10^{44}\rm erg/s$ which falls in the region of the upper limits placed on the X-ray luminosity of LRDs should they be AGN-dominated (Ananna et al., 2024). Given our population of LRDs is most comparable to the brightest observed LRDs, it is unsurprising that our LRDs have X-ray luminosities near the upper limits from observations. This would also allow for lower mass, dimmer LRDs with similar overall structure to those found in Astrid to fall below the upper limits placed on AGN $L_{X}$ from observational nondetection.
@@ 5 Where Are The Dimmer LRDs?
The population of LRDs in Astrid represents a subset of the observed population, as their number density is significantly lower than those found in observations, and they are concentrated within the brighter portion of the observed LRD population. Astrid contains $\sim 1/100$ of the expected number of LRDs with UV magnitudes brighter than $\rm M_{UV}=-19$, but completely lacks LRDs with $\rm M_{UV}\geq-18.5$ resulting in an overall LRD number density of $\sim 1/1000$ of observations (Kocevski et al., 2024; Kokorev et al., 2024; Pérez-González et al., 2024).
@@
This galaxy has an F277W-F [^4] W color of 0.88, which is insufficient to host an LRD. The primary differences between the spectrum of this galaxy and that of the example LRD in Figure 2 are the relative brightness of the AGN and the dust attenuation of both components. In the LRD spectrum, before dust attenuation, the AGN and stellar components have similar brightness in the F [^4] W band. Despite having one of the brightest AGN relative to its stellar mass among the lower stellar mass galaxies, the AGN emission of the dim LRD candidate galaxy is only $\sim 16\%$ as bright as its host galaxy in the F [^4] W band before dust attenuation. This indicates very few, if any, of the lower mass galaxies in Astrid have AGN that are bright enough to significantly impact their overall color. This is likely the result of both black hole accretion and star formation being induced by the presence of cold dense gas, producing host galaxies which are brighter due to their increased stellar mass, and young stellar population. This coupling is broken once AGN feedback begins quenching the host galaxy, as we see in our LRDs and models of high redshift galaxies hosting over-massive black holes (Pacucci and Loeb, 2024; Pacucci et al., 2024).
@@
These two factors are heavily impacted by modeling choices in our mock observation pipeline. AGN bolometric luminosity is determined by the simulation, but the AGN SED can impact how that luminosity is distributed, and is dependent on the properties of the gas very near the black hole, which is not resolved by cosmological simulations. Recent work (e.g., Naidu et al., 2025), suggests that AGN embedded in a denser gas cloud, similar to models of direct collapse black holes (e.g. Pacucci et al., 2015, 2016), may be a strong candidate for LRD emission, which produces a significantly different SED than our more traditional model. Similarly, our dust model has a normalization factor calibrated in Bird et al. (2022), but may not be tuned appropriately for LRD galaxies.
We experiment with these two modeling choices on our dim LRD candidate galaxy. We increase the hydrogen column density of our AGN SED model in cloudy from $10^{23}$ $\rm N/cm^{2}$ to $10^{24}$ $\rm N/cm^{2}$ to approximate the effect of the ”gas-enshrouded" AGN models proposed in works like Inayoshi and Maiolino (2025); Naidu et al. (2025) and, Taylor et al. (2025). Additionally, we increase the dust normalization by a factor of four (from $\kappa_{\rm ISM}=10^{4.1}$ to $\kappa_{\rm ISM}=10^{4.7}$) for the stellar emission to explore the impact of errors in normalization. This also serves as an approximate test of the impact of potential under-density due to simulation effects like resolution limitations and gravitational softening.
@@
The “gas-enshrouded" AGN SED introduces another reddening mechanism to these sources, which could also address the lack of very red LRDs in Astrid. Given the “gas-enshrouded" AGN model we adopted in this experiment was fairly conservative (Naidu et al. (2025) found a Hydrogen column density of $\rm N(H)/cm^{2}\sim 10^{26}$) it is likely that a more extreme version of this model would have a significantly redder spectrum. If such an SED were applied to Eddington or Super-Eddington sources, they may appear similar to the reddest objects seen in observations. We intend to explore these alternative AGN models in more depth in future work.
The increased hydrogen column density in the AGN SED introduces both a steeper intrinsic AGN SED and a significant Balmer break, both of which contribute to the red appearance of this galaxy. These qualities are especially important for any LRDs with minimal dust content, like those suggested by Setton et al. (2025); Xiao et al. (2025), and would result in the consistent Balmer break transition that is found in Setton et al. (2024).
The increased stellar dust attenuation has two primary effects: (I) it reddens the stellar spectrum and (II) it decreases the brightness of the stellar component, which increases the relative contribution of the AGN to the total emission of the source. This increase in AGN contribution is particularly impactful given the significant red color of the “gas-enshrouded" AGN SED. If these high Eddington black holes were present in dimmer, lower mass host galaxies, that would also produce sources with heightened AGN contributions. With the black hole models used in Astrid this is unlikely to occur due to the black hole seeding and feedback models, but it may be possible under alternative models. The BRAHMA simulation suite (Bhowmick et al., 2024) suggests that the over-massive black holes seen in observations are reproducible in simulation with specific seeding techniques, and that much of their mass is the result of black hole mergers. This would allow for the formation of more massive black holes in lower mass galaxies, as their mass growth would be less coupled to star formation. These more massive black holes would be capable of brighter AGN emission without exceeding the Eddington limit, potentially appearing as LRDs.